### v1

# Model Already Knows the Best Noise: Bayesian Active Noise Selection via Attention in Video Diffusion Model

Abstract
The choice of initial noise significantly affects the quality and prompt alignment of video diffusion models, where different noise seeds for the same prompt can lead to drastically different generations. While recent methods rely on externally designed priors such as frequency filters or inter-frame smoothing, they often overlook internal model signals that indicate which noise seeds are inherently preferable. To address this, we propose ANSE (Active Noise Selection for Generation), a model-aware framework that selects high-quality noise seeds by quantifying attention-based uncertainty. At its core is BANSA (Bayesian Active Noise Selection via Attention), an acquisition function that measures entropy disagreement across multiple stochastic attention samples to estimate model confidence and consistency. For efficient inference-time deployment, we introduce a Bernoulli-masked approximation of BANSA that enables score estimation using a single diffusion step and a subset of attention layers. Experiments on CogVideoX-2B and 5B demonstrate that ANSE improves video quality and temporal coherence with only an 8% and 13% increase in inference time, respectively, providing a principled and generalizable approach to noise selection in video diffusion. See our project page: https://anse-project.github.io/anse-project/
Figure 1: Random Seed vs. Ours. We propose ANSE, a noise selection framework, and the BANSA Score, an uncertainty-based metric. By selecting initial noise seeds with lower BANSA scores, which indicate more certain noise samples, ANSE improves video generation performance.
$\dagger$\$$\dagger$\$footnotetext: First and corresponding author
Figure 2: Conceptual comparison of noise initialization. (a) Prior methods (1; 2) iteratively refine noise using frequency domain priors through full diffusion sampling, incurring significant computational cost. (b) In contrast, our approach selects optimal noise seeds by estimating attention-based uncertainty at the first denoising step, enabling efficient and model-aware noise selection.
1 Introduction
Diffusion models have rapidly established themselves as a powerful class of generative models, demonstrating state-of-the-art performance across images and videos (3; 4; 5; 6; 7; 8; 9; 10; 11). In particular, Text-to-Video (T2V) diffusion models have received increasing attention for their ability to generate temporally coherent and visually rich video sequences. To achieve this, most T2V model architectures extend Text-to-Image (T2I) diffusion backbones by incorporating temporal modules or motion-aware attention layers (12; 13; 14; 15; 7; 16; 8). Furthermore, other works explore video generative structures, such as causal autoencoders or video autoencoder-based models, which aim to generate full video volumes rather than a sequence of independent frames (17; 9; 18; 19; 10; 11).
Beyond architectural design, another promising direction lies in improving noise initialization at inference time for T2I and T2V generation (20; 21; 22; 23). This aligns with the growing trend of inference-time scaling, observed not only in Large Language Models (24; 25) but also in diffusion-based generation systems (26). Due to the iterative nature of the diffusion process, the choice of initial noise profoundly influences video quality, temporal consistency, and prompt alignment (27; 1; 28; 2). As illustrated in Figure 1, the same prompt can lead to drastically different videos depending solely on the noise seed, motivating the need for intelligent noise selection.
Several recent approaches attempt to address this by designing external noise priors. For example, PYoCo (27) introduces inter-frame dependent noise patterns to improve coherence, though it requires extensive fine-tuning. FreeNoise (28) reschedules noise across time using a fusion-based strategy, while FreeInit (1) applies frequency-domain filtering to preserve low-frequency components. FreqPrior (2) extends this idea via Gaussian-shaped frequency priors and partial sampling. While effective, these methods rely on externally designed priors and require multiple full diffusion passes to evaluate candidate seeds. More importantly, they fail to leverage internal signals within the model that indicate which noise seeds are inherently preferable.
To address this limitation, we propose a model-aware noise selection framework, ANSE (Active Noise Selection for Generation), grounded in Bayesian uncertainty. At the core of ANSE is BANSA (Bayesian Active Noise Selection via Attention), an acquisition function that identifies noise seeds inducing confident and consistent attention behaviors under stochastic perturbations. A conceptual comparison between our method and prior frequency-based approaches is illustrated in Figure 2, highlighting the difference between external priors and model-informed uncertainty estimates.
Unlike BALD(29), which operates on classification logits, BANSA measures entropy in attention maps, arguably the most informative signals in generative diffusion. It compares the average entropy of individual maps to the entropy of their mean, capturing both uncertainty and disagreement across forward passes. A low BANSA score indicates that the model is more confident and certain in its attention, which empirically correlates with coherent video generation, as shown in Figure 1.
To make this approach suitable for inference-time, we approximate BANSA using Bernoulli-masked attention, which enables multiple stochastic attention samples from a single forward pass. We further reduce computation by limiting BANSA evaluation to early denoising steps and a subset of informative attention layers, selected via correlation analysis. Our contributions are threefold:
  • We present ANSE, the first active noise selection framework for video diffusion models, built on a principled Bayesian formulation of attention-based uncertainty.
  • We introduce BANSA, a novel acquisition function that quantifies attention consistency under stochastic perturbations, enabling model-aware noise selection without retraining or external noise priors.
  • Our method enhances both video quality and temporal consistency across various text-to-video architectures, with only a marginal increase in inference time of about 8% for CogVideoX-2B and 13% for CogVideoX-5B.
2 Preliminary
Video Diffusion Models
Diffusion models (30; 31) have achieved remarkable success across generative tasks. For T2V generation, directly operating in pixel space incurs high computational cost. To address this, video diffusion models (VDMs) typically adopt the latent diffusion model (LDM) framework, where the diffusion process is conducted in a compressed latent space.
A video autoencoder, composed of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, is trained to reconstruct the original input video $\mathbf{x}$ such that $\mathbf{x}=\mathcal{D}(\mathcal{E}(\mathbf{x}))$. Denoting the latent code as $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x})$, the forward diffusion process adds noise over time:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\quad t=1,\dots,T,$
where $\bar{\alpha}_{t}$ is a pre-defined variance schedule. To learn the reverse process, a denoising network $\boldsymbol{\epsilon}_{\theta}$ is trained using the denoising score matching loss (32):
$\mathcal{L}_{\theta}=\mathbb{E}_{\mathbf{z}_{t},\boldsymbol{\epsilon},t}\left[\left\|\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{c},t)-\boldsymbol{\epsilon}\right\|^{2}\right],$
where $\mathbf{c}$ denotes the conditioning text prompt. During sampling, the generation begins from Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ and proceeds via a deterministic DDIM solver (33). The update at each step is computed as:
$\mathbf{z}_{t-1}=\sqrt{\bar{\alpha}_{t}}\,\hat{\mathbf{z}}_{0}(t)+\sqrt{1-\bar{\alpha}_{t-1}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{c},t),$
where the denoised latent estimate $\hat{\mathbf{z}}_{0}(t):=\frac{\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{c},t)}{\sqrt{\bar{\alpha}_{t}}}$ is obtained using Tweedie’s formula (34; 35). This iterative process continues until $t=1$, yielding the final denoised latent $\mathbf{z}_{0}$, which is decoded into a video via $\mathcal{D}$.
Bayesian Active Learning by Disagreement (BALD)
Active Learning improves model performance by selecting the most informative samples from an unlabeled pool in training phase. Acquisition functions are typically categorized into uncertainty-based (36; 29; 37; 38) and distribution-based (39; 40; 41; 42) approaches, with some relying on external modules such as auxiliary predictors (38; 43; 44). While active learning has been predominantly applied to image classification tasks, in this work, we focus on adapting uncertainty-based methods to text-to-video generation, without requiring additional models.
Predictive entropy is a common uncertainty measure, but it captures only aleatoric uncertainty and fails to account for parameter uncertainty. BALD addresses this by quantifying epistemic uncertainty via the mutual information between predictions $\mathbf{y}$ and model parameters $\theta$:
$\mathrm{BALD}(\mathbf{x})=\mathcal{H}[p(\mathbf{y}|\mathbf{x})]-\mathbb{E}_{p(\theta|\mathcal{D}_{U})}\left[\mathcal{H}[p(\mathbf{y}|\mathbf{x},\theta)]\right],$ (1)
where $\mathcal{H}[p]=-\sum_{y}p(y)\log p(y)$ is the Shannon entropy (45). The first term captures the entropy of the mean prediction, while the second term averages the entropy over stochastic forward passes. A high BALD score indicates confident but disagreeing predictions, revealing high epistemic uncertainty.
Since the posterior over $\theta$ is intractable, BALD is approximated using $K$ stochastic forward passes (e.g., Monte Carlo dropout):
$\widehat{\mathrm{BALD}}(\mathbf{x})=\mathcal{H}\left[\frac{1}{K}\sum_{k=1}^{K}p^{(k)}(\mathbf{y}|\mathbf{x})\right]-\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}\left[p^{(k)}(\mathbf{y}|\mathbf{x})\right].$ (2)
We reinterpret BALD for inference-time generative modeling. Rather than selecting samples for labeling, we apply BALD to rank noise seeds by their epistemic uncertainty. Selecting seeds with lower BALD scores results in more stable model behavior and leads to higher-quality generations.
Figure 3: Overview of our BANSA-based noise selection process. Given a text prompt $c$, we compute BANSA scores for multiple noise seeds $\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$ using Bernoulli-masked attention maps from selected layers at an early diffusion step. The seed with the lowest score, indicating confident and consistent attention, is selected for generation.
3 Methods
We propose ANSE, a framework for selecting high-quality noise seeds in T2V diffusion models based on model uncertainty as shown in Figure 2. ANSE is built upon an acquisition function called BANSA, which extends uncertainty-based criteria from classification tasks to the attention space of generative diffusion models (Section 3.1). To enable efficient inference-time application, we approximate BANSA using Bernoulli-masked attention sampling (Section 3.2). Furthermore, to reduce computational redundancy, we identify a representative attention layer using correlation-based linear probing (Section 3.3). The overall pipeline is illustrated in Figure 3.
3.1 BANSA: Bayesian Active Noise Selection via Attention
We introduce BANSA, an acquisition function for selecting optimal noise seeds in T2V diffusion models. Unlike classification tasks with explicit predictive distributions, diffusion models lack such outputs. We instead estimate uncertainty in the attention space, where alignment between text and visual tokens naturally emerges during generation. Here, attention maps are treated as stochastic predictions conditioned on the noise seed $\mathbf{z}$, prompt $c$, and diffusion timestep $t$. BANSA measures disagreement and confidence across multiple attention samples, capturing attention-level uncertainty analogous to BALD, but tailored to the generative setting.
Definition 1 (BANSA Score).
Let $\mathbf{z}$ be a noise seed, $c$ a text prompt, and $t$ a diffusion timestep. Let $\mathbf{Q}(\mathbf{z},c,t),\mathbf{K}(\mathbf{z},c,t)\in\mathbb{R}^{N\times d}$ denote the query and key matrices from a denoising network $\epsilon_{\theta}$. The attention map is computed as:
$A(\mathbf{z},c,t):=\texttt{Softmax}\left(\mathbf{Q}(\mathbf{z},c,t)\,\mathbf{K}(\mathbf{z},c,t)^{\top}\right)\in\mathbb{R}^{N\times N}.$ (3)
Let $\mathcal{A}(\mathbf{z},c,t)=\{A^{(1)},\dots,A^{(K)}\}$ denote a set of $K$ stochastic attention maps obtained via forward passes with random perturbations (e.g., Bernoulli masking). The BANSA score is defined as:
$\text{BANSA}(\mathbf{z},c,t):=\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)})-\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right),$ (4)
where the row-wise entropy of an attention map $A$ is given by:
$\mathcal{H}(A):=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}-A_{ij}\log A_{ij}.$ (5)
This formulation captures both the sharpness (confidence) and the consistency (agreement) of attention behavior. BANSA can be applied to various attention types (e.g., cross-, self-, or temporal) and allows layer-wise interpretability.
Given a noise pool $\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$, we select the optimal noise seed that minimizes the BANSA score:
$\mathbf{z}^{*}:=\arg\min_{\mathbf{z}\in\mathcal{Z}}\text{BANSA}(\mathbf{z},c,t).$ (6)
A desirable property of BANSA is that its score becomes zero when all attention samples are identical, reflecting complete agreement and certainty. We formalize this as follows:
Proposition 1 (BANSA Zero Condition).
Let $\mathcal{A}(\mathbf{z},c,t)=\{A^{(1)},\dots,A^{(K)}\}$ be a set of row-stochastic attention maps. Then:
$\text{BANSA}(\mathbf{z},c,t)=0\quad\Leftrightarrow\quad A^{(1)}=\cdots=A^{(K)}.$
The proof is deferred to the Appendix. This condition implies that minimizing the BANSA score promotes attention behavior that is both confident and consistent under stochastic perturbations. Empirically, such attention patterns are associated with better prompt alignment, temporal coherence, and visual fidelity in generated videos. Therefore, BANSA serves as a principled criterion for model-aware noise selection in T2V.
3.2 Stochastic Approximation of BANSA via Bernoulli-Masked Attention
While BANSA provides a principled objective for noise selection, its computation requires $K$ independent forward passes per noise seed $\mathbf{z}$, which is computationally expensive. To mitigate this cost, we propose a stochastic approximation using Bernoulli-masked attention, enabling multiple attention samples from a single pass.
Instead of computing stochastic attention maps from $K$ separate forward passes such as equipping dropout, we inject stochasticity directly into the attention computation by applying binary masks to the attention scores. For each sample iteration $k=1,\cdots,K$, we generate a binary mask $m_{k}\in\{0,1\}^{N\times N}$ where each element is drawn i.i.d. from $\text{Bernoulli}(p)$. The masked attention map is computed as:
$\hat{A}^{(k)}(\mathbf{z},c,t):=\texttt{Softmax}\left(\left(\mathbf{Q}(\mathbf{z},c,t)\,\mathbf{K}(\mathbf{z},c,t)^{\top}\right)\odot m_{k}\right).$ (7)
where $\odot$ denotes element-wise multiplication. These masks simulate variability in attention patterns while keeping the input $(\mathbf{z},c,t)$ fixed. Using $K$ such samples, we define the approximate BANSA:
[…]
Input: Text prompt $c$, noise pool $\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$, timestep $t$, cutoff layer $d^{*}$
Output: Generated video $\hat{v}$
1 foreach $\mathbf{z}_{i}\in\mathcal{Z}$ do
2       Compute BANSA score: ${\widehat{\text{BANSA-E}}}_{\leq d^{*}}(\mathbf{z}_{i},c,t)$ via Eq. (9);
3
4Select optimal noise: $\mathbf{z}^{*}=\arg\min_{\mathbf{z}_{i}}{\widehat{\text{BANSA-E}}}_{\leq d^{*}}(\mathbf{z}_{i},c,t)$;
5 Generate video: $\hat{v}=\texttt{SampleVideo}(\mathbf{z}^{*},c,t)$;
return $\hat{v}$
Algorithm 1 Active Noise Selection with BANSA Score for Video Generation
3.3 Layer Selection via Cumulative BANSA Correlation
BANSA can be computed at any attention layer, but attention behavior varies across depth. While using all layers provides a comprehensive uncertainty estimate, it is computationally heavy for deep T2V models. To address this, we propose a correlation-based truncation strategy that selects the smallest depth $d^{*}$ such that the averaged BANSA score over the first $d$ layers remains highly correlated with the full-layer score.
[…]
${\widehat{\text{BANSA-E}}}_{\leq d}(\mathbf{z}_{i},c,t):=\frac{1}{d}\sum_{l=1}^{d}{\text{BANSA-E}}^{(l)}(\mathbf{z}_{i},c,t).\vspace{-0.5em}$ (9)
To determine $d^{*}$, we compute the Pearson correlation (46) between ${\widehat{\text{BANSA-E}}}_{\leq d}$ and the full-layer average ${\widehat{\text{BANSA-E}}}_{\leq L}$, and select the smallest $d$ satisfying:
$\text{Corr}\left({\widehat{\text{BANSA-E}}}_{\leq d},{\widehat{\text{BANSA-E}}}_{\leq L}\right)\geq\tau,$ (10)
with $\tau=0.7$ in our experiments. We validate this procedure using 100 prompts and 10 noise seeds across CogVideoX-2B, and CogVideoX-5B. As shown in Figure 5, the correlation stabilizes at layer 14 in CogVideoX-2B, and 19 in CogVideoX-5B. We therefore set $d^{*}$ accordingly and define the BANSA score as ${\widehat{\text{BANSA-E}}}_{\leq d^{*}}$ to guide noise selection, as summarized in Algorithm 1.
This layer selection procedure provides a lightweight and model-specific approximation of the full BANSA score. Since $d^{*}$ can be predefined for each model, it does not interfere with the noise sampling process and introduces no runtime cost during generation. As shown in the Appendix, ${\widehat{\text{BANSA-E}}}_{\leq d^{*}}$ closely approximates the full-layer score and achieves comparable generation quality across all models.
4 Experiments
Experimental Setting. We evaluate ANSE on two representative text-to-video (T2V) diffusion models: CogVideoX-2B and CogVideoX-5B (9), chosen for their strong spatiotemporal modeling capabilities grounded in real-world dynamics. This setup enables rigorous evaluation of noise selection where attention and coherence are critical. We follow the official DDIM sampling protocol with 50 denoising steps for both models. Quantitative results for noise prior based approaches such as Freeinit (1) and FreqPrior (2) are omitted, as they are not officially supported on CogVideoX and incur 3$\times$ inference cost. Nonetheless, ANSE is orthogonal and can be combined with these methods for further gains. We use a noise pool of size $M{=}10$ with $K{=}10$ stochastic forward passes per noise, and apply Bernoulli-masked attention with a masking probability $p{=}0.2$. Additional details are in the Appendix. All experiments are run on NVIDIA H100 GPUs.
Evaluation Metric. To evaluate the impact of ANSE, we use VBench (47), a perceptually grounded benchmark for text-to-video generation. VBench reports two high-level metrics—quality score and semantic score—which are combined into a total score via weighted averaging and normalized to a 0–100 scale. Each score is a composite metric: the quality score is derived from 7 perceptual dimensions including subject consistency, background consistency, temporal flickering, motion smoothness, dynamic degree, aesthetic quality, and imaging quality; the semantic score comprises 9 alignment-related criteria such as object class, multiple objects, human action, color, spatial relationship, scene, temporal style, appearance style, and overall consistency. This decomposition ensures that VBench comprehensively assesses both visual fidelity and semantic alignment. For each configuration (with and without BANSA), we generate 4,730 videos to ensure statistical reliability.
Table 1: Quantitative results on VBench using CogVideoX-2B and -5B. ANSE consistently improves quality, semantic alignment, and total score. Scores for noise prior methods (1; 2) are omitted as they are not officially supported on CogVideoX and require costly re-implementation (3$\times$ inference), making large-scale evaluation impractical.
Model Backbone Method Quality Score Semantic Score Total Score Inference Time
CogVideoX-2B (9) Vanilla 82.08 76.83 81.03 247.8
+ Ours 82.56(+0.48) 78.06(+1.23) 81.66(+0.63) 269.3(+8.68%)
CogVideoX-5B (9) Vanilla 82.53 77.50 81.52 1223.5
+ Ours 82.70(+0.17) 78.10(+0.60) 81.71(+0.25) 1392.1(+13.78%)
Figure 4: Qualitative comparison of CogVideoX variants with and without ANSE. Results from CogVideoX-2B are shown in the first two rows; the rest show results from CogVideoX-5B. With ANSE, videos exhibit improved visual quality, better text alignment, and smoother motion transitions compared to the baseline.
Quantitative Comparison. As shown in Table 1, ANSE consistently improves performance across both CogVideoX models. On CogVideoX-2B, the total VBench score increases from 81.03 (Vanilla) to 81.66 with ANSE, driven by gains in quality (+0.48) and semantic alignment (+1.23). On the larger CogVideoX-5B, ANSE also improves all metrics: quality increases from 82.53 to 82.70 (+0.17), semantic score from 77.50 to 78.10 (+0.60), and total score from 81.52 to 81.72 (+0.25). These results demonstrate that ANSE effectively enhances both perceptual fidelity and semantic alignment, even on large-scale, temporally grounded video diffusion models. The consistent gains across both model scales highlight the robustness and generalizability of ANSE in selecting high-quality noise seeds under varying architectural complexities.
Qualitative Comparison. Figure 4 shows qualitative results on CogVideoX-2B and -5B with and without ANSE. Our method improves semantic fidelity, motion portrayal, and visual clarity across diverse prompts. For example, in "exploding" and ANSE captures key semantic transitions—generating visible explosions and preserving temporal continuity. In "vacuuming", the subject remains static in the baseline but exhibits purposeful motion with ANSE.
On CogVideoX-5B, similar improvements are evident. In "koala playing the piano" and "zebra running", ANSE generates anatomically coherent bodies with expressive motion. These results demonstrate ANSE’s ability to enhance spatial-temporal fidelity and generalize to high-capacity video diffusion models.
Computational Cost. As shown in Table 1, ANSE increases inference time by only +8% on CogVideoX-2B and +13% on CogVideoX-5B, measured in denoising steps. This overhead stems from noise seed evaluation but does not affect the sampling process, so memory usage remains unchanged. In contrast, prior methods such as Freeinit and FreqPrior require three full sampling passes, resulting in a 200% increase in inference time. While these methods have not been officially implemented, making direct comparisons challenging, ANSE reduces inference cost by approximately 64% while achieving comparable or superior generation quality.
5 Ablation Study and Analysis
Table 2: Comparison of different acquisition functions for noise selection.
Method Quality Score Semantic Score Total Score
Random 82.08 76.83 81.03
Entropy 82.23 76.73 81.13
BANSA (D) 82.43 76.91 81.33
BANSA (B) 82.56 78.06 81.66
Table 3: Effect of varying the number of $K$.
$K$ Subject Consistency Background Consistnecy
1 0.9618 0.9788
3 0.9623 0.9793
5 0.9632 0.9798
7 0.9638 0.9802
10 0.9641 0.9811
Figure 5: Correlation analysis between cumulative BANSA score and full-layer scores. The 0.7 threshold is reached around layer 14 for CogVideoX-2B and layer 19 for CogVideoX-5B.
Figure 6: Ablation study on noise pool size $M$. We evaluate total scores across three text-to-video diffusion models with varying $M$, and select suitable values based on computational cost.
Table 4: Quantitative comparison of reversed BANSA scoring on CogVideoX-2B. This presents results when selecting samples using the highest BANSA scores, compared to the default selection.
Method Subject Consistency Background Consistency Temporal Flickering Motion Smoothness Aesthetic Quality Imaging Quality Dynamic Degree Quality Score
Vanilla 0.9616 0.9788 0.9715 0.9743 0.6195 0.6267 0.6380 82.08
+ Ours (reverse) 0.9626 0.9785 0.9700 0.9741 0.6181 0.6253 0.6328 81.93
+ Ours 0.9641 0.9811 0.9775 0.9746 0.6202 0.6276 0.6511 82.56
Comparison of Acquisition Functions. We compare BANSA with alternative acquisition strategies for noise seed selection using the CogVideoX-2B model. As shown in Table 3, we evaluate random sampling, entropy-based selection, and two BANSA variants: BANSA (B), which uses Bernoulli masking, and BANSA (D), which introduces Dropout-based stochasticity. While all methods improve over the baseline, BANSA (B) consistently achieves the highest scores across quality, semantic, and total metrics. This confirms that injecting uncertainty through Bernoulli masking is more effective than dropout, which is commonly used in Bayesian acquisition functions such as BALD. The result highlights the importance of modeling attention-level uncertainty in a manner that reflects the structure of the underlying model.
Effect of Ensemble Size $K$. We assess the impact of the number of stochastic forward passes $K$ on subject and background consistency, again using CogVideoX-2B. As shown in Table 3, both metrics improve steadily as $K$ increases from 1 to 10, suggesting that larger ensembles yield more robust and stable noise evaluations. Performance saturates at $K=10$, which we adopt as the default throughout all experiments.
Effect of Noise Pool Size $M$. We analyze the role of noise pool size $M$, which determines the diversity of candidate seeds assessed by BANSA. While a larger $M$ increases the likelihood of discovering high-quality seeds, it also raises inference cost. As shown in Figure 6, performance saturates around $M=10$ for CogVideoX-2B and -5B. We set these values as defaults for each model.
Reversing the BANSA Criterion. To further validate BANSA, we conduct a control experiment where the noise seed with the highest BANSA score is selected—i.e., choosing the seed associated with the greatest model uncertainty. As shown in Table 4, this reversal results in degradation of quality-related metrics, confirming that lower BANSA scores are predictive of perceptually stronger generations and supporting the validity of our selection strategy.
Figure 7: Failure case and limitation of our method. Although the BANSA score indicates low uncertainty, the resulting video still contains unnatural content. This represent a limitation of ours: we select optimal seeds but do not alter the generation process itself.
6 Discussion and Limitations
Our method focuses on noise seed selection through model uncertainty estimation, yet it has notable limitations. As shown in Figure 7, even seeds with low BANSA scores, which indicate high model confidence, can produce unnatural generations. This suggests that while ANSE effectively identifies promising initial seeds, it does not directly affect the generation process itself. Moreover, there remains a gap between estimated uncertainty and perceptual quality. Although BANSA reliably captures attention-level uncertainty, it may not fully reflect semantic or aesthetic aspects. While generating multiple candidates per seed and selecting based on strong quality metrics would be ideal, this approach is computationally prohibitive. We thus view BANSA as a practical surrogate for such strategies. Future work could further enhance performance by integrating it with information-theoretic refinement or active learning methods.
7 Conclusion
We present ANSE, a framework for active noise selection in video diffusion models. It is built around BANSA, an acquisition function that uses attention-derived uncertainty to identify noise seeds that promote confident and consistent attention, which are indicative of high-quality generation. BANSA adapts the BALD principle to the generative setting by operating in the attention space. To enable efficient deployment, we introduce a stochastic approximation using Bernoulli-masked attention and a lightweight layer selection method. Experiments across multiple T2V backbones show that ANSE improves video quality and prompt alignment with little to no increase in inference time. This work highlights the potential of inference-time noise selection guided by internal model signals. By extending active learning concepts beyond training-time data selection, ANSE enables model-aware decisions during inference and enhances generation without retraining. Our approach introduces a new inference-time scaling paradigm, where performance is improved not by modifying the model or increasing sampling steps, but through informed seed selection.
Appendix A Supplementary Section
In this supplementary document, we present the following:
  • Implementation details of the BANSA score in Section B.
  • Proof of Proposition 1 from the main paper regarding the BANSA Zero condition in Section C.
  • Implementation details of ANSE and evaluation metrics in Section D.
  • Further explanation of layer selection through cumulative BANSA correlation in Section E.
  • Additional ablation studies including full-layer BANSA Score analysis in Section F and temporal scope effects in Section G.
  • Additional qualitative results demonstrating the impact of BANSA Score in Section H.
Appendix B Implementation Details of BANSA Score
Score Definition vs. Implementation.
In the main paper, we define the BANSA score as the difference between the mean of entropies and the entropy of the mean:
$\text{BANSA}(\mathbf{z},c,t):=\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)})-\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right),$
This corresponds to the negative of the mutual information formulation used in BALD (29). But, for practical and interpretability reasons, we adopt the BALD-style computation in our implementation:
$\text{BANSA}(\mathbf{z},c,t):=\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right)-\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)}),$
which yields a non-negative score. This implementation allows easier visualization and interpretation, where larger values correspond to higher disagreement (i.e., uncertainty) across stochastic attention maps. We adopt this sign convention to align with standard BALD implementations, where mutual information is expressed as a non-negative measure of epistemic uncertainty. All figures and tables in the paper (e.g., correlation plots and performance curves) reflect this convention, showing positive-valued BANSA scores.
Semantic consistency.
Although the mathematical sign differs between the definition and implementation, the semantic meaning and selection behavior are strictly preserved. In both cases, we minimize the BANSA score to select noise seeds with low attention uncertainty. The relative ordering of scores remains unchanged, and the interpretation of the score as a measure of model disagreement is fully retained.
For clarity, we note that the proof of Proposition 1 in the following section is also presented under this BALD-style convention.
Appendix C Proof of Proposition 1
Proposition 1 (BANSA Zero Condition).
Let $\mathcal{A}(\mathbf{z},c,t)=\{A^{(1)},\dots,A^{(K)}\}$ be a set of row-stochastic attention maps. Then:
$\text{BANSA}(\mathbf{z},c,t)=0\quad\Leftrightarrow\quad A^{(1)}=\cdots=A^{(K)}.$
Proof.
BANSA is defined as the difference between the average entropy and the entropy of the average:
$\text{BANSA}(\mathbf{z},c,t)=\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}(\mathbf{z},c,t)\right)-\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)}(\mathbf{z},c,t)).$
Since the Shannon entropy $\mathcal{H}(\cdot)$ is strictly concave over the probability simplex. Therefore, by Jensen’s inequality:
$\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right)\geq\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)}),$
with equality if and only if $A^{(1)}=\cdots=A^{(K)}$. Thus, $\text{BANSA}(\mathbf{z},c,t)=0$ if and only if all attention maps are identical. ∎
Remark 1.
(Interpretation) This result confirms that the BANSA score quantifies disagreement among sampled attention maps. A BANSA score of zero occurs only when all stochastic attention realizations collapse to a single deterministic map—i.e., the model exhibits no epistemic uncertainty in its attention distribution. Higher BANSA values indicate greater variation across samples, and thus, higher uncertainty. In this sense, BANSA acts as a Jensen–Shannon-type divergence over attention maps, capturing their dispersion under stochastic masking.
Appendix D Further Details on Evaluation Metrics and Implementation
Evaluation Metrics
To evaluate performance on Vbench, we use the Vbench-long version, where prompts are augmented using GPT-4o across all evaluation dimensions. This version is specifically designed for assessing videos longer than 4 seconds.
We rigorously evaluate our generated videos following the official evaluation protocol. The Quality Score is a weighted average of the following aspects: subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, imaging quality, and dynamic degree.
The Semantic Score is a weighted average of the following semantic dimensions: object class, multiple objects, human action, color, spatial relationship, scene, appearance style, temporal style, and overall consistency.
The Total Score is then computed as a weighted combination of the Quality Score and Semantic Score:
Total $\text{Total Score}=\frac{w_{1}}{w_{1}+w_{2}}\cdot\text{Quality Score}+\frac{w_{2}}{w_{1}+w_{2}}\cdot\text{Semantic Score}$
where $w_{1}=4$ and $w_{2}=1$, following the default setting in the official implementation.
Implementation
As discussed in Section B, we compute our BANSA score using the BALD-style formulation, which yields non-negative values. For clearer visualization, we normalize the BANSA scores from their original range (minimum: 0.45, maximum: 0.60) to the [0, 1] interval. This normalization is used solely for visual clarity in figures and plots, and does not affect the noise selection process, which operates on the raw BANSA scores.
Appendix E Further Detail of BANSA Layer-wise Correlation Analysis
Prompt construction.
We evenly sampled 100 prompts from the four official VBench categories: Subject Consistency, Overall Consistency, Temporal Flickering, and Scene. Each category contains 25 prompts, selected to ensure diversity in motion, structure, and semantics.
Below are representative examples:
  • Subject Consistency (e.g., “A young man with long, flowing hair sits on a rustic wooden stool in a cozy room, strumming an acoustic guitar…”)
  • Overall Consistency (e.g., “A mesmerizing splash of turquoise water erupts in extreme slow motion, each droplet suspended in mid-air…”)
  • Temporal Flickering (e.g., “A cozy restaurant with flickering candles and soft music. Patrons dine peacefully as snow falls outside…”)
  • Scene (e.g., “A university campus transitions from lively student life to a golden sunset behind the clock tower…”)
Prompt sampling was stratified to ensure coverage of diverse visual and temporal patterns. The full list of prompts will be made publicly available upon code release.
BANSA score computation and correlation analysis.
For each prompt, we generated 10 videos using different random noise seeds and computed BANSA scores at each attention layer. This yielded one full-layer BANSA score and a set of layer-wise scores per seed.
To obtain stable estimates, we averaged the per-layer and full-layer BANSA scores across the 10 seeds, reducing noise-specific variance and capturing consistent uncertainty patterns.
We then computed Pearson correlations between the cumulative BANSA scores (summed from layer 1 to $d$) and the official quality scores. The optimal depth $d^{*}$ was defined as the smallest $d$ at which the correlation exceeded 0.7. This procedure is visualized in Figure 5 of the main paper, and $d^{*}$ was applied consistently throughout all experiments. This setup ensures that our correlation analysis reflects generalizable, noise-agnostic trends in attention-based uncertainty.
Table 5: Comparison between full-layer and truncated BANSA score.
Backbone Model Method Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Temporal Flickering $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Dynamic Degree $\uparrow$ Quality Score $\uparrow$ Inference Time $\downarrow$
CogvideoX-2B Full-layer 0.9639 0.9810 0.9801 0.9743 0.6198 0.6244 0.6516 82.58 303.7
Truncated 0.9641 0.9811 0.9775 0.9746 0.6202 0.6276 0.6511 82.56 269.3
CogvideoX-5B Full-layer 0.9660 0.9630 0.9863 0.9708 0.6168 0.6290 0.6979 82.71 1530.1
Truncated 0.9658 0̱.9639 0.9861 0.9711 0.6179 0.6290 0.6918 82.70 1392.3
Appendix F Effectiveness of Truncated BANSA Score
To reduce the computational overhead of BANSA evaluation, we adopt a truncated score that aggregates attention uncertainty only up to a fixed depth $d^{*}$, rather than summing over all layers. To evaluate the effectiveness of this approximation, we compared the final generation quality when selecting noise seeds using either the full-layer or truncated BANSA scores.
As shown in Table 5, both approaches yield highly similar results across all seven dimensions of the VBench evaluation protocol (subject consistency, background consistency, aesthetic quality, imaging quality, motion smoothness, dynamic degree, and temporal flickering). Importantly, the overall quality scores are preserved despite the substantial reduction in attention layers used.
This demonstrates that truncated BANSA is sufficient to capture the key uncertainty signals for reliable noise selection while reducing inference time. The strong alignment in quality stems from the fact that our method relies on relative ranking rather than absolute values, allowing for efficient yet robust selection with significantly lower computational cost. We attribute this effectiveness to the fact that most informative attention behaviors emerge early in the denoising process, allowing accurate uncertainty estimation without full-layer computation.
Table 6: Effect of temporal scope in BANSA score on generation quality.
BANSA Scope Subject Consistency $\uparrow$ Temporal Flickering $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Dynamic Degree $\uparrow$ Inference Time $\downarrow$
1-step 0.9639 0.9801 0.9743 0.6198 0.6244 0.6516 $\times$ 1
25-step avg 0.9651 0.9798 0.9746 0.6202 0.6271 0.6511 $\times$ 25
50-step avg 0.9652 0.9799 0.9751 0.6203 0.6276 0.6514 $\times$ 50
Appendix G Effect of Temporal Scope in BANSA Score
While our method computes the BANSA score only at the first denoising step to minimize cost, it is natural to ask whether incorporating more timesteps improves its predictive power for noise selection. To investigate this, we compute the average BANSA score across the first 1, 25, and 50 denoising steps and compare their effectiveness in predicting video quality.
Table 6 reports the VBench scores for subject consistency, aesthetic quality, imaging quality, motion smoothness, dynamic degree, and temporal flickering when using BANSA computed over different temporal scopes. Although using more timesteps results in slightly better quality, the gains are marginal. This indicates that most of the predictive signal for noise quality is embedded early in the generation trajectory.
More importantly, since BANSA is used solely to assess the uncertainty of the initial noise seed—not to track full-step generation behavior—our 1-step computation is sufficient to capture the core uncertainty signal. In contrast, computing BANSA over all steps requires running multiple attention forward passes across the full trajectory, resulting in substantial computational overhead that limits its practicality for real-world applications.
Appendix H Additional Qualitative Comparison
More qualitative results.
Figures 8 and 9 present additional examples generated using our noise selection framework. Across diverse prompts, the selected seeds yield improved spatial detail, aesthetic quality, and semantic alignment, further validating the robustness of our approach. These examples complement our quantitative findings by illustrating the visual impact of BANSA-based noise selection.
Effect of BANSA score on generation quality.
Figures 10 provide a qualitative comparison of outputs generated using three types of noise seeds: a randomly sampled seed, the seed with the highest BANSA score (lowest quality), and the seed with the lowest BANSA score (highest quality). All videos were generated using 50 denoising steps with the CogVideoX-5B backbone. The lowest-BANSA seed consistently produces sharper, more coherent, and semantically faithful videos, whereas the highest-BANSA seed often leads to structural artifacts or temporal instability. These results highlight the practical value of BANSA-guided noise selection.
Figure 8: Effect of ANSE on semantic fidelity and motion stability in CogVideoX outputs. Each block compares baseline generations with those using ANSE-selected noise. Across both CogVideoX-2B and 5B, ANSE improves semantic alignment to the prompt and reduces artifacts such as temporal flickering and object distortion.
Figure 9: Additional qualitative comparison of CogVideoX variants with and without ANSE. Results from CogVideoX-2B are shown in the first two rows; the rest show CogVideoX-5B. With ANSE, videos exhibit improved visual quality, better text alignment, and smoother motion transitions compared to the baseline.
Figure 10: Qualitative comparison of generations from different noise seeds. We compare outputs generated from a randomly sampled seed (top), the seed with the highest BANSA score (middle), and the seed with the lowest score (bottom), using the same prompt and model. BANSA-selected seeds produce more coherent structure, stable motion, and stronger semantic alignment than both random and high-uncertainty seeds.